can you haelp me solve the system of equations. using the Gaussian elimination qith back substitution or gauss-jordan elimination
x+5y= 0
x+6y+z=1
-2x-y-z=-41
can you break it down to me on how to solve problems of this type

Answer by vertciel(183)

Hello,
This is a system of three equations. You have to eliminate for only each term, and then you use this term to help you solve the others.
x + 5y = 0 ----------(1)
x + 6y + z = 1 ------(2)
-2x - y - z = -41 ---(3)
(2) + (3): -x + 5y = -40 ---(4) <-- I am doing this so I can get rid of the z for the time being.
(4) + (1): 10y = -40
y = -4
Since you know that x + 5y = 0, x = -5y.
Therefore, x = -5(-4)
x = 20
I'll leave it to you to find z. Please write back for more help if needed.